\begin{thebibliography}{40}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Bapna and Firat(2019)}]{bapna-firat-2019-simple}
Ankur Bapna and Orhan Firat. 2019.
\newblock \href {https://doi.org/10.18653/v1/D19-1165} {Simple, scalable
  adaptation for neural machine translation}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 1538--1548, Hong Kong,
  China. Association for Computational Linguistics.

\bibitem[{Briakou et~al.(2021)Briakou, Lu, Zhang, and
  Tetreault}]{briakou2021xformal}
Eleftheria Briakou, Di~Lu, Ke~Zhang, and Joel~R. Tetreault. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.256} {Ol{\'{a}},
  bonjour, salve! {XFORMAL:} {A} benchmark for multilingual formality style
  transfer}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {NAACL-HLT} 2021, Online, June 6-11, 2021}, pages 3199--3216.
  Association for Computational Linguistics.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'{a}}n, Grave, Ott, Zettlemoyer, and
  Stoyanov}]{conneau2020unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzm{\'{a}}n, Edouard Grave, Myle Ott, Luke Zettlemoyer,
  and Veselin Stoyanov. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.747} {Unsupervised
  cross-lingual representation learning at scale}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2020, Online, July 5-10, 2020}, pages
  8440--8451. Association for Computational Linguistics.

\bibitem[{Conneau and Lample(2019)}]{lample2019crosslingual}
Alexis Conneau and Guillaume Lample. 2019.
\newblock \href
  {https://proceedings.neurips.cc/paper/2019/hash/c04c19c2c2474dbf5f7ac4372c5b9af1-Abstract.html}
  {Cross-lingual language model pretraining}.
\newblock In \emph{Advances in Neural Information Processing Systems 32: Annual
  Conference on Neural Information Processing Systems 2019, NeurIPS 2019,
  December 8-14, 2019, Vancouver, BC, Canada}, pages 7057--7067.

\bibitem[{Dale et~al.(2021)Dale, Voronov, Dementieva, Logacheva, Kozlova,
  Semenov, and Panchenko}]{dale2021text}
David Dale, Anton Voronov, Daryna Dementieva, Varvara Logacheva, Olga Kozlova,
  Nikita Semenov, and Alexander Panchenko. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.emnlp-main.629} {Text
  detoxification using large pre-trained neural models}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing, {EMNLP} 2021, Virtual Event / Punta Cana,
  Dominican Republic, 7-11 November, 2021}, pages 7979--7996. Association for
  Computational Linguistics.

\bibitem[{Dementieva et~al.(2021{\natexlab{a}})Dementieva, Moskovskiy,
  Logacheva, Dale, Kozlova, Semenov, and Panchenko}]{dementieva2021methods}
Daryna Dementieva, Daniil Moskovskiy, Varvara Logacheva, David Dale, Olga
  Kozlova, Nikita Semenov, and Alexander Panchenko. 2021{\natexlab{a}}.
\newblock \href {https://doi.org/10.3390/mti5090054} {Methods for
  detoxification of texts for the russian language}.
\newblock \emph{Multimodal Technol. Interact.}, 5(9):54.

\bibitem[{Dementieva et~al.(2021{\natexlab{b}})Dementieva, Ustyantsev, Dale,
  Kozlova, Semenov, Panchenko, and Logacheva}]{dementieva2021crowdsourcing}
Daryna Dementieva, Sergey Ustyantsev, David Dale, Olga Kozlova, Nikita Semenov,
  Alexander Panchenko, and Varvara Logacheva. 2021{\natexlab{b}}.
\newblock \href {http://ceur-ws.org/Vol-2932/paper2.pdf} {Crowdsourcing of
  parallel corpora: the case of style transfer for detoxification}.
\newblock In \emph{Proceedings of the 2nd Crowd Science Workshop: Trust,
  Ethics, and Excellence in Crowdsourced Data Management at Scale co-located
  with 47th International Conference on Very Large Data Bases (VLDB 2021
  (https://vldb.org/2021/))}, pages 35--49, Copenhagen, Denmark. CEUR Workshop
  Proceedings.

\bibitem[{Devaraj et~al.(2021)Devaraj, Marshall, Wallace, and
  Li}]{devaraj2021paragraphlevel}
Ashwin Devaraj, Iain Marshall, Byron Wallace, and Junyi~Jessy Li. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.395}
  {Paragraph-level simplification of medical texts}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 4972--4984, Online. Association for Computational
  Linguistics.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin2019bert}
Jacob Devlin, Ming{-}Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/n19-1423} {{BERT:} pre-training of
  deep bidirectional transformers for language understanding}.
\newblock In \emph{Proceedings of the 2019 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {NAACL-HLT} 2019, Minneapolis, MN, USA, June 2-7, 2019, Volume
  1 (Long and Short Papers)}, pages 4171--4186. Association for Computational
  Linguistics.

\bibitem[{dos Santos et~al.(2018{\natexlab{a}})dos Santos, Melnyk, and
  Padhi}]{santos2018fighting}
Cicero~Nogueira dos Santos, Igor Melnyk, and Inkit Padhi. 2018{\natexlab{a}}.
\newblock \href {http://arxiv.org/abs/1805.07685} {Fighting offensive language
  on social media with unsupervised text style transfer}.

\bibitem[{dos Santos et~al.(2018{\natexlab{b}})dos Santos, Melnyk, and
  Padhi}]{nogueira-dos-santos-etal-2018-fighting}
C{\'{\i}}cero~Nogueira dos Santos, Igor Melnyk, and Inkit Padhi.
  2018{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/P18-2031} {Fighting offensive
  language on social media with unsupervised text style transfer}.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2018, Melbourne, Australia, July 15-20,
  2018, Volume 2: Short Papers}, pages 189--194. Association for Computational
  Linguistics.

\bibitem[{Feng et~al.(2020)Feng, Yang, Cer, Arivazhagan, and
  Wang}]{feng2020languageagnostic}
Fangxiaoyu Feng, Yinfei Yang, Daniel Cer, Naveen Arivazhagan, and Wei Wang.
  2020.
\newblock \href {http://arxiv.org/abs/2007.01852} {Language-agnostic {BERT}
  sentence embedding}.
\newblock \emph{CoRR}, abs/2007.01852.

\bibitem[{Houlsby et~al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone,
  de~Laroussilhe, Gesmundo, Attariyan, and
  Gelly}]{DBLP:conf/icml/HoulsbyGJMLGAG19}
Neil Houlsby, Andrei Giurgiu, Stanislaw Jastrzebski, Bruna Morrone, Quentin
  de~Laroussilhe, Andrea Gesmundo, Mona Attariyan, and Sylvain Gelly. 2019.
\newblock \href {http://proceedings.mlr.press/v97/houlsby19a.html}
  {Parameter-efficient transfer learning for {NLP}}.
\newblock In \emph{Proceedings of the 36th International Conference on Machine
  Learning, {ICML} 2019, 9-15 June 2019, Long Beach, California, {USA}},
  volume~97 of \emph{Proceedings of Machine Learning Research}, pages
  2790--2799. {PMLR}.

\bibitem[{Jigsaw(2018)}]{jigsaw_toxic}
Jigsaw. 2018.
\newblock Toxic comment classification challenge.
\newblock
  \href{https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge}{https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge}.
\newblock Accessed: 2021-03-01.

\bibitem[{Jin et~al.(2020)Jin, Jin, Hu, Vechtomova, and Mihalcea}]{jin2021deep}
Di~Jin, Zhijing Jin, Zhiting Hu, Olga Vechtomova, and Rada Mihalcea. 2020.
\newblock \href {http://arxiv.org/abs/2011.00416} {Deep learning for text style
  transfer: {A} survey}.
\newblock \emph{CoRR}, abs/2011.00416.

\bibitem[{Kingma and Ba(2015)}]{kingma2017adam}
Diederik~P. Kingma and Jimmy Ba. 2015.
\newblock \href {http://arxiv.org/abs/1412.6980} {Adam: {A} method for
  stochastic optimization}.
\newblock In \emph{3rd International Conference on Learning Representations,
  {ICLR} 2015, San Diego, CA, USA, May 7-9, 2015, Conference Track
  Proceedings}.

\bibitem[{Krishna et~al.(2020)Krishna, Wieting, and
  Iyyer}]{krishna-etal-2020-reformulating}
Kalpesh Krishna, John Wieting, and Mohit Iyyer. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.55} {Reformulating
  unsupervised style transfer as paraphrase generation}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 737--762, Online. Association for
  Computational Linguistics.

\bibitem[{Lai et~al.(2022)Lai, Toral, and Nissim}]{multilingual_tst}
Huiyuan Lai, Antonio Toral, and Malvina Nissim. 2022.
\newblock \href {https://doi.org/10.48550/arXiv.2203.08552} {Multilingual
  pre-training with language and task adaptation for multilingual text style
  transfer}.
\newblock \emph{CoRR}, abs/2203.08552.

\bibitem[{Laugier et~al.(2021)Laugier, Pavlopoulos, Sorensen, and
  Dixon}]{laugier2021civil}
Leo Laugier, John Pavlopoulos, Jeffrey Sorensen, and Lucas Dixon. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.eacl-main.124} {Civil
  rephrases of toxic texts with self-supervised transformers}.
\newblock In \emph{Proceedings of the 16th Conference of the European Chapter
  of the Association for Computational Linguistics: Main Volume, {EACL} 2021,
  Online, April 19 - 23, 2021}, pages 1442--1461. Association for Computational
  Linguistics.

\bibitem[{Lewis et~al.(2020)Lewis, Liu, Goyal, Ghazvininejad, Mohamed, Levy,
  Stoyanov, and Zettlemoyer}]{lewis2019bart}
Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed,
  Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.703} {{BART:}
  denoising sequence-to-sequence pre-training for natural language generation,
  translation, and comprehension}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2020, Online, July 5-10, 2020}, pages
  7871--7880. Association for Computational Linguistics.

\bibitem[{Li et~al.(2021)Li, Chen, Yang, Gao, Zhao, and
  Yan}]{li2020stylecontent}
Mingzhe Li, Xiuying Chen, Min Yang, Shen Gao, Dongyan Zhao, and Rui Yan. 2021.
\newblock \href {https://ojs.aaai.org/index.php/AAAI/article/view/17565} {The
  style-content duality of attractiveness: Learning to write eye-catching
  headlines via disentanglement}.
\newblock In \emph{Thirty-Fifth {AAAI} Conference on Artificial Intelligence,
  {AAAI} 2021, Thirty-Third Conference on Innovative Applications of Artificial
  Intelligence, {IAAI} 2021, The Eleventh Symposium on Educational Advances in
  Artificial Intelligence, {EAAI} 2021, Virtual Event, February 2-9, 2021},
  pages 13252--13260. {AAAI} Press.

\bibitem[{Liu et~al.(2020)Liu, Gu, Goyal, Li, Edunov, Ghazvininejad, Lewis, and
  Zettlemoyer}]{liu2020multilingual}
Yinhan Liu, Jiatao Gu, Naman Goyal, Xian Li, Sergey Edunov, Marjan
  Ghazvininejad, Mike Lewis, and Luke Zettlemoyer. 2020.
\newblock \href {https://transacl.org/ojs/index.php/tacl/article/view/2107}
  {Multilingual denoising pre-training for neural machine translation}.
\newblock \emph{Trans. Assoc. Comput. Linguistics}, 8:726--742.

\bibitem[{Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis,
  Zettlemoyer, and Stoyanov}]{liu2019roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019.
\newblock \href {http://arxiv.org/abs/1907.11692} {Roberta: {A} robustly
  optimized {BERT} pretraining approach}.
\newblock \emph{CoRR}, abs/1907.11692.

\bibitem[{Maddela et~al.(2021)Maddela, Alva{-}Manchego, and
  Xu}]{maddela2021controllable}
Mounica Maddela, Fernando Alva{-}Manchego, and Wei Xu. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.277} {Controllable
  text simplification with explicit paraphrasing}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {NAACL-HLT} 2021, Online, June 6-11, 2021}, pages 3536--3553.
  Association for Computational Linguistics.

\bibitem[{Prabhumoye et~al.(2018)Prabhumoye, Tsvetkov, Black, and
  Salakhutdinov}]{prabhumoye2018style}
Shrimai Prabhumoye, Yulia Tsvetkov, Alan~W. Black, and Ruslan Salakhutdinov.
  2018.
\newblock \href {http://arxiv.org/abs/1809.06284} {Style transfer through
  multilingual and feedback-based back-translation}.
\newblock \emph{CoRR}, abs/1809.06284.

\bibitem[{Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena,
  Zhou, Li, and Liu}]{raffel2020exploring}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael
  Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu. 2020.
\newblock \href {http://jmlr.org/papers/v21/20-074.html} {Exploring the limits
  of transfer learning with a unified text-to-text transformer}.
\newblock \emph{J. Mach. Learn. Res.}, 21:140:1--140:67.

\bibitem[{Rao and Tetreault(2018)}]{rao-tetreault-2018-dear}
Sudha Rao and Joel Tetreault. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-1012} {Dear sir or madam, may
  {I} introduce the {GYAFC} dataset: Corpus, benchmarks and metrics for
  formality style transfer}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 129--140, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{Sharma et~al.(2021)Sharma, Lin, Miner, Atkins, and
  Althoff}]{sharma2021facilitating}
Ashish Sharma, Inna~W. Lin, Adam~S. Miner, David~C. Atkins, and Tim Althoff.
  2021.
\newblock \href {https://doi.org/10.1145/3442381.3450097} {Towards facilitating
  empathic conversations in online mental health support: {A} reinforcement
  learning approach}.
\newblock In \emph{{WWW} '21: The Web Conference 2021, Virtual Event /
  Ljubljana, Slovenia, April 19-23, 2021}, pages 194--205. {ACM} / {IW3C2}.

\bibitem[{Shen et~al.(2017)Shen, Lei, Barzilay, and Jaakkola}]{shen2017style}
Tianxiao Shen, Tao Lei, Regina Barzilay, and Tommi~S. Jaakkola. 2017.
\newblock \href
  {https://proceedings.neurips.cc/paper/2017/hash/2d2c8394e31101a261abf1784302bf75-Abstract.html}
  {Style transfer from non-parallel text by cross-alignment}.
\newblock In \emph{Advances in Neural Information Processing Systems 30: Annual
  Conference on Neural Information Processing Systems 2017, December 4-9, 2017,
  Long Beach, CA, {USA}}, pages 6830--6841.

\bibitem[{Tan et~al.(2019)Tan, Ren, He, Qin, Zhao, and
  Liu}]{tan2019multilingual}
Xu~Tan, Yi~Ren, Di~He, Tao Qin, Zhou Zhao, and Tie{-}Yan Liu. 2019.
\newblock \href {https://openreview.net/forum?id=S1gUsoR9YX} {Multilingual
  neural machine translation with knowledge distillation}.
\newblock In \emph{7th International Conference on Learning Representations,
  {ICLR} 2019, New Orleans, LA, USA, May 6-9, 2019}. OpenReview.net.

\bibitem[{Wang et~al.(2019{\natexlab{a}})Wang, Hua, and
  Wan}]{wang2019controllable}
Ke~Wang, Hang Hua, and Xiaojun Wan. 2019{\natexlab{a}}.
\newblock \href
  {https://proceedings.neurips.cc/paper/2019/hash/8804f94e16ba5b680e239a554a08f7d2-Abstract.html}
  {Controllable unsupervised text attribute transfer via editing entangled
  latent representation}.
\newblock In \emph{Advances in Neural Information Processing Systems 32: Annual
  Conference on Neural Information Processing Systems 2019, NeurIPS 2019,
  December 8-14, 2019, Vancouver, BC, Canada}, pages 11034--11044.

\bibitem[{Wang et~al.(2019{\natexlab{b}})Wang, Wu, Mou, Li, and
  Chao}]{Wang2019HarnessingPN}
Yunli Wang, Yu~Wu, Lili Mou, Zhoujun Li, and Wenhan Chao. 2019{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/D19-1365} {Harnessing pre-trained
  neural networks with rules for formality style transfer}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing, {EMNLP-IJCNLP} 2019, Hong Kong, China, November
  3-7, 2019}, pages 3571--3576. Association for Computational Linguistics.

\bibitem[{Warstadt et~al.(2019)Warstadt, Singh, and
  Bowman}]{warstadt2019neural}
Alex Warstadt, Amanpreet Singh, and Samuel~R. Bowman. 2019.
\newblock \href {https://transacl.org/ojs/index.php/tacl/article/view/1710}
  {Neural network acceptability judgments}.
\newblock \emph{Trans. Assoc. Comput. Linguistics}, 7:625--641.

\bibitem[{Wieting et~al.(2019)Wieting, Berg-Kirkpatrick, Gimpel, and
  Neubig}]{wieting2019beyond}
John Wieting, Taylor Berg-Kirkpatrick, Kevin Gimpel, and Graham Neubig. 2019.
\newblock \href {https://arxiv.org/abs/1909.06694} {Beyond bleu: Training
  neural machine translation with semantic similarity}.
\newblock In \emph{Proceedings of the Association for Computational
  Linguistics}.

\bibitem[{Wieting and Gimpel(2018)}]{wieting-gimpel-2018-paranmt}
John Wieting and Kevin Gimpel. 2018.
\newblock \href {https://doi.org/10.18653/v1/P18-1042} {{P}ara{NMT}-50{M}:
  Pushing the limits of paraphrastic sentence embeddings with millions of
  machine translations}.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 451--462,
  Melbourne, Australia. Association for Computational Linguistics.

\bibitem[{Wu et~al.(2019)Wu, Lv, Zang, Han, and Hu}]{condbert}
Xing Wu, Shangwen Lv, Liangjun Zang, Jizhong Han, and Songlin Hu. 2019.
\newblock Conditional bert contextual augmentation.
\newblock In \emph{Computational Science -- ICCS 2019}, pages 84--95, Cham.
  Springer International Publishing.

\bibitem[{Xu et~al.(2021)Xu, Lu, Sun, Ma, and Guo}]{xu2021vae}
Haoran Xu, Sixing Lu, Zhongkai Sun, Chengyuan Ma, and Chenlei Guo. 2021.
\newblock \href {http://arxiv.org/abs/2112.03154} {{VAE} based text style
  transfer with pivot words enhancement learning}.
\newblock \emph{CoRR}, abs/2112.03154.

\bibitem[{Xue et~al.(2021)Xue, Constant, Roberts, Kale, Al{-}Rfou, Siddhant,
  Barua, and Raffel}]{xue2021mt5}
Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al{-}Rfou, Aditya
  Siddhant, Aditya Barua, and Colin Raffel. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.41} {mt5: {A}
  massively multilingual pre-trained text-to-text transformer}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {NAACL-HLT} 2021, Online, June 6-11, 2021}, pages 483--498.
  Association for Computational Linguistics.

\bibitem[{Yao and Yu(2021)}]{yao2021improving}
Zonghai Yao and Hong Yu. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.124} {Improving
  formality style transfer with context-aware rule injection}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing, {ACL/IJCNLP} 2021, (Volume 1: Long Papers),
  Virtual Event, August 1-6, 2021}, pages 1561--1570. Association for
  Computational Linguistics.

\bibitem[{Zhang et~al.(2020)Zhang, Ge, and Sun}]{zhang2020parallel}
Yi~Zhang, Tao Ge, and Xu~Sun. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.294} {Parallel data
  augmentation for formality style transfer}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2020, Online, July 5-10, 2020}, pages
  3221--3228. Association for Computational Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{acl}

\usepackage{times}
\usepackage{latexsym}
\usepackage[russian,english]{babel}
\usepackage[multiple]{footmisc}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\title{Exploring Cross-lingual Textual Style Transfer with \\ Large Multilingual Language Models}

  \author{Daniil Moskovskiy$^1$~~~~~~~ Daryna Dementieva$^{1,2}$~~~~~~~ Alexander Panchenko$^1$ \\
  $^1$Skolkovo Institute of Science and Technology, Russia\\
  $^2$Technical University of Munich, Germany\\
  \texttt{\href{mailto:daniil.moskovskiy@skoltech.ru}{\{daniil.moskovskiy,daryna.dementieva,a.panchenko\}@skoltech.ru}} }

\date{}

\begin{document}
\maketitle
\begin{abstract}
Detoxification is a task of generating text in polite style while preserving meaning and fluency of the original toxic text. Existing detoxification methods are designed to work in one exact language. This work investigates  multilingual and cross-lingual detoxification and the behavior of large multilingual models like in this setting. Unlike previous works we aim to make large language models able to perform detoxification without direct fine-tuning in given language. Experiments show that multilingual models are capable of performing multilingual style transfer. However, models are not able to perform cross-lingual detoxification and direct fine-tuning on exact language is inevitable. 
\end{abstract}

\section{Introduction}

The task of Textual Style Transfer (Textual Style Transfer) can be viewed as a task where certain properties of text are being modified while rest retain the same\footnote{Hereinafter the data-driven definition of style is used. Therefore, we call style a characteristic of given dataset that differs from a general dataset \cite{jin2021deep}.}. In this work we focus on detoxification textual style transfer \cite{santos2018fighting,dementieva2021methods}. It can be formulated as follows: given two text corpora $D^X=\{x_1, x_2, \dots x_n\}$ and $D^Y=\{y_1, y_2, \dots, y_n\}$, where $X$, $Y$ - are two sets of all possible text in styles $s^X$, $s^Y$ respectively, we want to build a model $f_{\theta} : X \rightarrow Y$, such that the probability $p(y_{gen}|x, s^X, s^Y)$ of transferring the style $s^X$ of given text $x$ (by generation $y_{gen}$) to the style $s^Y$ is maximized (where $s^X$ and $s^Y$ are toxic and non-toxic styles respectively).

Some examples of detoxification presented in Table \ref{tab:detoxification_example}.

Textual style transfer gained a lot of attention with a rise of deep learning-based NLP methods. Given that, Textual Style Transfer has now a lot of specific subtasks ranging from formality style transfer \cite{rao-tetreault-2018-dear, yao2021improving} and simplification of domain-specific texts \cite{devaraj2021paragraphlevel, maddela2021controllable} to emotion modification \cite{sharma2021facilitating} and detoxification (debiasing) \cite{li2020stylecontent, dementieva2021methods}. 

\begin{table*}[]
\centering
\begin{tabular}{c|c}
\hline
\textbf{Source text}         & \textbf{Target text}  \\ \hline
What the f*ck is your problem?  & What is your problem?     \\
This whole article is bullshit. & This article is not good. \\
Yeah, this clowns gonna make alberta great again! & Yeah, this gonna make Alberta great again \\ \hline
\end{tabular}
\caption{Examples of desired detoxification results.}
\label{tab:detoxification_example}
\end{table*}

There exist a variety of Textual Style Transfer methods: from totally \textbf{supervised} methods \cite{Wang2019HarnessingPN,zhang2020parallel,dementieva2021methods} which require a parallel text corpus for training to \textbf{unsupervised} \cite{shen2017style,wang2019controllable,xu2021vae} that are designed to work without any parallel data. The latter sub-field of research is more popular nowadays due to the scarcity of parallel text data for Textual Style Transfer. On the other hand, if we address Textual Style Transfer task as a Machine Translation task we get a significant performance boost \cite{prabhumoye2018style}. 

The task of detoxification, in which we focus in this work, is relatively new. First work on detoxification was a sequence-to-sequence collaborative classifier, attention and the cycle consistency loss \cite{nogueira-dos-santos-etal-2018-fighting}. A recent work by \cite{laugier2021civil} introduces self-supervised model based on T$5$ model \cite{raffel2020exploring}  with a denoising and cyclic auto-encoder loss. 

Both these methods are unsupervised which is an advantage but it comes from the major current problem of the textual style transfer. There is a lack of parallel data for Textual Style Transfer since there exist only few parallel datasets for English \cite{rao-tetreault-2018-dear} and some other languages \cite{briakou2021xformal}. When it comes to detoxification there are only two parallel detoxification corpora available now and they both appeared only last year \cite{dementieva2021crowdsourcing}. Most state-of-the-art methods rely on large amounts of text data which is often available for some well-researched languages like English but lacking for other languages almost entirely. Therefore, it is important to study whether cross-lingual (or at least multilingual) detoxification is possible.



Multilingual language models such as mBART \cite{liu2020multilingual}, mT5 \cite{xue2021mt5} have recently become available. This work explores the possibility of multilingual and cross-lingual textual style transfer (Textual Style Transfer) using such large multilingual language models. We test the hypothesis that modern large text-to-text models are able to generalize ability of style transfer across languages. 

Our contributions can be summarized as follows\footnote{All code is available online: \url{https://github.com/skoltech-nlp/multilingual_detox}}:
\begin{enumerate}
    \item We introduce a novel study of multilingual textual style transfer and conduct experiments with several multilingual language models and evaluate their performance. 
    \item We conduct cross-lingual Textual Style Transfer experiments to investigate whether multilingual language models are able to perform Textual Style Transfer without fine-tuning on a specific language.
\end{enumerate}

\section{Methodology}

We formulate the task of \textbf{supervised} Textual Style Transfer as a sequence-to-sequence NMT task and fine-tune multilingual language models to translate from "toxic" to "polite" language.

\subsection{Datasets}

In this work we use two datasets for Russian and English languages. Aggregated information about datasets could be found in Table \ref{table:kysymys}, examples from datasets can be found in \ref{sec:appendix_1} and \ref{sec:appendix_2}.

\begin{table}[h!]
\centering
\label{tab:table_1}
\begin{tabular}{l| l l l}
\hline
\textbf{Language}        & \textbf{Train}    & \textbf{Dev}    & \textbf{Test}   \\ \hline
English & $18 777$ & $988$  & $671$  \\ 
Russian & $5058$   & $1000$ & $1000$ \\
\hline
\end{tabular}
\caption{Aggregated datasets statistics.}
\label{table:kysymys}
\end{table}

\paragraph{Russian data} We use detoxification dataset\footnote{\url{https://github.com/skoltech-nlp/russe_detox_2022}} which consists of $5 058$ training sentences, $1 000$ validation sentences and $1 000$ test sentences.

\paragraph{English data} We use ParaDetox \cite{dementieva2021crowdsourcing} dataset. It consists of $19766$ \textit{toxic} sentences and their \textit{polite} paraphrases. This data is split into training and validation as $95\%$ for training and $5\%$ for validation. For testing we use a set of $671$ toxic sentences. 

\subsection{Experimental Setup}

We perform a series of experiments on detoxification using parallel data for English and Russian. We train models in two different setups: \textbf{multilingual} and \textbf{cross-lingual}. 

\paragraph{Multilingual setup} In this setup we train models on data containing both English and Russian texts
and then compare their performance with baselines trained on these languages solely.

\paragraph{Cross-lingual setup} In cross-lingual setup we test the hypothesis that models are able to perform detoxification without explicit fine-tuning on exact language. We fine-tune models on English and Russian separately and then test their performance.   

\subsection{Models}

Scaling language models to many languages has become an emerging topic of interest recently \cite{devlin2019bert, tan2019multilingual, lample2019crosslingual, conneau2020unsupervised}. We adopt several multilingual models to textual style transfer in our work. 


\paragraph{Baselines} We use two detoxification methods as baselines in this work - \textbf{Delete} method which simply deletes toxic words in the sentence according to the vocabulary of toxic words and \textbf{CondBERT}. The latter approach works in usual masked-LM setup by masking toxic words and replacing them with non-toxic ones. This approach was first proposed by \cite{condbert} as a data augmentation method and then adopted to detoxification by \cite{dale2021text}.  

\paragraph{mT5} mT5 \cite{xue2021mt5} is a multilingual version of T5 \cite{raffel2020exploring} - a text-to-text transformer model, which was trained on many downstream tasks. mT5 replicates T5 training but now it is trained on more than $100$ languages.

\paragraph{mBART} mBART \cite{liu2020multilingual} is a multilingual variation of BART  \cite{lewis2019bart} - denoising autoencoder built with a sequence-to-sequence model. mBART is trained on monolingual corpora across many languages. We adopt mBART in sequence-to-sequence detoxification task via fine-tuning on parallel detoxification dataset.  



\subsection{Evaluation metrics}

Unlike other NLP tasks, one metric is not enough to benchmark the quality of style transfer. The ideal Textual Style Transfer model output should \textit{preserve the original content} of the text, \textit{change the style} of the original text to target and the generated text also should \textit{be grammatically correct}. We follow \newcite{dale2021text} approach in Textual Style Transfer evaluation.


\subsubsection{Content Preservation}

\paragraph{Russian} Content preservation score (\textbf{SIM}) is evaluated as a cosine similarity of LaBSE \cite{feng2020languageagnostic} sentence embeddings. The model is slightly different from the original one, only English and Russian embeddings are left. 

\paragraph{English} Similarity (\textbf{SIM}) between the embedding of the original sentence and the generated one is calculated using the model presented by \newcite{wieting2019beyond}. Being is trained on paraphrase pairs extracted from ParaNMT corpus \cite{wieting-gimpel-2018-paranmt}, this model's training objective is to select embeddings such that the similarity of embeddings of paraphrases is higher than the similarity between sentences that are not paraphrases.

\subsubsection{Grammatic and language quality (fluency)}

\paragraph{Russian} We measure fluency (\textbf{FL}) with a BERT-based classifier \cite{devlin2019bert}
trained to distinguish real texts from corrupted ones. The model was trained on Russian texts and their corrupted (random word replacement, word deletion and insertion, word shuffling etc.) versions. Fluency is calculated as a difference between the probabilities of being corrupted for source and target sentences. The logic behind using difference is that we ensure that the generated sentence is not worse than the original one in terms of fluency.

\paragraph{English} We measure fluency (\textbf{FL}) as a percentage of fluent sentences evaluated by the RoBERTa-based\footnote{\url{https://huggingface.co/roberta-large}} \cite{liu2019roberta} classifier of linguistic acceptability trained on CoLA \cite{warstadt2019neural} dataset.

\subsubsection{Style transfer accuracy}

\paragraph{Russian} Style transfer accuracy (\textbf{STA}) is evaluated with a BERT-based \cite{devlin2019bert} toxicity classifier\footnote{\url{https://huggingface.co/SkolkovoInstitute/russian_toxicity_classifier}} fine-tuned from RuBERT Conversational. This classifier was additionally trained on Russian Language Toxic Comments dataset collected from \url{2ch.hk} and Toxic Russian Comments dataset collected from \url{ok.ru}. 

\paragraph{English} Style transfer accuracy (\textbf{STA}) is calculated with a style classifier - RoBERTa-based \cite{liu2019roberta} model trained on the union of three Jigsaw datasets \cite{jigsaw_toxic}. The sentence is considered toxic when the classifier confidence is above $0.8$. The classifier reaches the AUC-ROC of $0.98$ and F$_1$-score of $0.76$.


\subsubsection{Joint metric}

Aforementioned metrics must be properly combined to get one \textit{Joint} metric to evaluate Textual Style Transfer. We follow \newcite{krishna-etal-2020-reformulating} and calculate \textbf{J} as an average of products    of sentence-level \textit{fluency}, \textit{style transfer accuracy}, and \textit{content
preservation}: 
\begin{equation}
    \textbf{J} = \frac{1}{n}\sum\limits_{i=1}^{n}\textbf{STA}(x_i) \cdot \textbf{SIM}(x_i) \cdot \textbf{FL}(x_i) 
\end{equation}


\begin{table*}[!htbp]
\footnotesize
\centering
\begin{tabular}{lllllllll}
\hline
\multicolumn{1}{|l|}{} &
  \multicolumn{1}{c|}{\textbf{STA}$\uparrow$} &
  \multicolumn{1}{c|}{\textbf{SIM}$\uparrow$} &
  \multicolumn{1}{c|}{\textbf{FL}$\uparrow$} &
  \multicolumn{1}{c|}{\textbf{J}$\uparrow$} &
  \multicolumn{1}{c|}{\textbf{STA}$\uparrow$} &
  \multicolumn{1}{c|}{\textbf{SIM}$\uparrow$} &
  \multicolumn{1}{c|}{\textbf{FL}$\uparrow$} &
  \multicolumn{1}{c|}{\textbf{J}$\uparrow$} \\ \hline
\multicolumn{1}{|c|}{} &
  \multicolumn{4}{c|}{\textbf{Russian}} &
  \multicolumn{4}{c|}{\textbf{English}} \\ \hline  
\multicolumn{1}{|c|}{} &
  \multicolumn{8}{c|}{\it Baselines} \\ \hline
\multicolumn{1}{|l|}{Delete} &
  \multicolumn{1}{l|}{0.532} &
  \multicolumn{1}{l|}{0.875} &
  \multicolumn{1}{l|}{0.834} &
  \multicolumn{1}{l|}{0.364} &
  \multicolumn{1}{l|}{0.810} &
  \multicolumn{1}{l|}{0.930} &
  \multicolumn{1}{l|}{0.640} &
  \multicolumn{1}{l|}{0.460} \\ 
\multicolumn{1}{|l|}{condBERT~\cite{dale2021text} }  &
  \multicolumn{1}{l|}{{0.819}} &
  \multicolumn{1}{l|}{0.778} &
  \multicolumn{1}{l|}{0.744} &
  \multicolumn{1}{l|}{0.422} &
  \multicolumn{1}{l|}{\textbf{0.980}} &
  \multicolumn{1}{l|}{0.770} &
  \multicolumn{1}{l|}{0.820} &
  \multicolumn{1}{l|}{0.620} \\ \hline
\multicolumn{1}{|c|}{} &
  \multicolumn{8}{c|}{\it  Multilingual Setup} \\ \hline
\multicolumn{1}{|l|}{mT$5$ base} &
      \multicolumn{1}{l|}{0.772} &
  \multicolumn{1}{l|}{0.676} &
  \multicolumn{1}{l|}{0.795} &
  \multicolumn{1}{l|}{0.430} &
  \multicolumn{1}{l|}{0.833} &
  \multicolumn{1}{l|}{0.826} &
  \multicolumn{1}{l|}{0.830} &
  \multicolumn{1}{l|}{0.556} \\
\multicolumn{1}{|l|}{mT$5$ small} &
  \multicolumn{1}{l|}{0.745} &
  \multicolumn{1}{l|}{0.705} &
  \multicolumn{1}{l|}{0.794} &
  \multicolumn{1}{l|}{0.428} &
  \multicolumn{1}{l|}{0.826} &
  \multicolumn{1}{l|}{0.841} &
  \multicolumn{1}{l|}{0.763} &
  \multicolumn{1}{l|}{0.513} \\
\multicolumn{1}{|l|}{mT$5$ base$^{*}$} &
  \multicolumn{1}{l|}{{0.773}} &
  \multicolumn{1}{l|}{0.676} &
  \multicolumn{1}{l|}{0.795} &
  \multicolumn{1}{l|}{0.430} &
  \multicolumn{1}{l|}{0.893} &
  \multicolumn{1}{l|}{0.787} &
  \multicolumn{1}{l|}{\textbf{0.942}} &
  \multicolumn{1}{l|}{\textbf{0.657}} \\
\multicolumn{1}{|l|}{mBART $5000$} &
  \multicolumn{1}{l|}{0.685} &
  \multicolumn{1}{l|}{\textbf{0.778}} &
  \multicolumn{1}{l|}{\textbf{0.841}} &
  \multicolumn{1}{l|}{\textbf{0.449}} &
  \multicolumn{1}{l|}{0.887} &
  \multicolumn{1}{l|}{\textbf{0.889}} &
  \multicolumn{1}{l|}{0.866} &
  \multicolumn{1}{l|}{0.640} \\ \hline
\multicolumn{1}{|c|}{} &
  \multicolumn{8}{c|}{\it  Cross-lingual Setup} \\ \hline
\multicolumn{1}{|l|}{mT$5$ base ENG} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.838}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.276}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.506}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.115}} &
  \multicolumn{1}{l|}{0.860} &
  \multicolumn{1}{l|}{0.834} &
  \multicolumn{1}{l|}{0.833} &
  \multicolumn{1}{l|}{0.587} \\
\multicolumn{1}{|l|}{mT$5$ base RUS} &
  \multicolumn{1}{l|}{0.676} &
  \multicolumn{1}{l|}{0.794} &
  \multicolumn{1}{l|}{0.846} &
  \multicolumn{1}{l|}{0.454} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.906}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.365}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.696}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.171}} \\
\multicolumn{1}{|l|}{mT$5$ small ENG} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.805}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.225}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.430}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.077}} &
  \multicolumn{1}{l|}{0.844} &
  \multicolumn{1}{l|}{0.858} &
  \multicolumn{1}{l|}{0.826} &
  \multicolumn{1}{l|}{0.591} \\
\multicolumn{1}{|l|}{mT$5$ small RUS} &
  \multicolumn{1}{l|}{0.559} &
  \multicolumn{1}{l|}{0.822} &
  \multicolumn{1}{l|}{0.817} &
  \multicolumn{1}{l|}{0.363} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.776}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.521}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.535}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.169}} \\
\multicolumn{1}{|l|}{mBART $3000$ ENG} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.923}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.395}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.552}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.202}} &
  \multicolumn{1}{l|}{0.842} &
  \multicolumn{1}{l|}{0.856} &
  \multicolumn{1}{l|}{0.876} &
  \multicolumn{1}{l|}{0.617} \\
\multicolumn{1}{|l|}{mBART $3000$ RUS} &
  \multicolumn{1}{l|}{0.699} &
  \multicolumn{1}{l|}{0.778} &
  \multicolumn{1}{l|}{0.858} &
  \multicolumn{1}{l|}{0.475} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.547}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.778}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.888}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.299}} \\
\multicolumn{1}{|l|}{mBART $5000$ ENG} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.900}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.299}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.591}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.160}} &
  \multicolumn{1}{l|}{0.857} &
  \multicolumn{1}{l|}{0.840} &
  \multicolumn{1}{l|}{0.873} &
  \multicolumn{1}{l|}{0.616} \\
\multicolumn{1}{|l|}{mBART $5000$ RUS} &
  \multicolumn{1}{l|}{0.724} &
  \multicolumn{1}{l|}{0.746} &
  \multicolumn{1}{l|}{0.827} &
  \multicolumn{1}{l|}{0.457} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.806}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.484}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.864}} &
  \multicolumn{1}{l|}{\textcolor{gray}{0.242}} \\ \hline

\multicolumn{1}{|c|}{} &
\multicolumn{8}{c|}{\it Backtranslation Setup} \\ \hline

\multicolumn{1}{|l|}{mBART $5000$ (Google)}  &
  \multicolumn{1}{l|}{0.675} &
  \multicolumn{1}{l|}{0.669} &
  \multicolumn{1}{l|}{0.634} &
  \multicolumn{1}{l|}{0.284} &
 \multicolumn{1}{l|}{0.678} &
 \multicolumn{1}{l|}{0.762} &
 \multicolumn{1}{l|}{0.568} &
\multicolumn{1}{l|}{0.284} \\ 
  
\multicolumn{1}{|l|}{mBART $5000$ (FSMT)} &
\multicolumn{1}{l|}{0.737} &
\multicolumn{1}{l|}{0.633} &
\multicolumn{1}{l|}{0.731} &
\multicolumn{1}{l|}{0.348} &
 \multicolumn{1}{l|}{0.744} &
 \multicolumn{1}{l|}{0.746} &
 \multicolumn{1}{l|}{0.893} &
 \multicolumn{1}{l|}{0.415} \\ \hline
\end{tabular}
\caption{Evaluation of TST models.
Numbers in \textbf{bold} indicate the best results. $\uparrow$ describes the higher the better metric. Results of unsuccessful TST depicted as \textcolor{gray}{gray}. ENG and RUS depicts the data model have been trained on. mT$5$ base$^{*}$ was trained on all English and Russian data available (datasets were not equalized). Last row depicts backtranslation workaround for cross-lingual detoxification. We include only the best result for brevity.} 
\label{tab:table_2}
\end{table*}

\subsection{Training}

There is a variety of versions of large multilingual models available. In this work we use small and base versions of mT5\footnote{\url{https://huggingface.co/google/mt5-base}}\footnote{\url{https://huggingface.co/google/mt5-large}} \cite{xue2021mt5} and large version of mBART\footnote{\url{https://huggingface.co/facebook/mbart-large-50-many-to-many-mmt}} \cite{liu2020multilingual}.

\paragraph{Multilingual training}

In multilingual training setup we fine-tune models using both English and Russian data. We use Adam \cite{kingma2017adam} optimizer for fine-tuning with different learning rates ranging from $1 \cdot 10^{-3}$ to $5 \cdot 10^{-5}$ with linear learning rate scheduling. We also test different number of warmup steps from $0$ to $1000$. We equalize Russian and English data for training and use $10000$ toxic sentences and their polite paraphrases for multilingual training in total. We train mT5 models for $40$ thousand  iterations\footnote{According to \cite{xue2021mt5} mT5 was not fine-tuned on downstream tasks as the original T5 model. Therefore, model requires more fine-tuning iterations for Textual Style Transfer.} with a batch size of $8$. We fine-tune mBART \cite{liu2020multilingual} for $1000$, $3000$, $5000$ and $10000$ iterations with batch size of $8$. 

\paragraph{Cross-lingual training}

In cross-lingual training setup we fine-tune models using only one dataset, e.g.: we fine-tune model on English data and check performance on both English and Russian data. Fine-tuning procedure was left the same: $40000$ iterations for mT5 models and $1000$, $3000$, $5000$ and $10000$ iterations for the mBART. 

\textbf{Back-translation approach} to cross-lingual style transfer proved to work substantially better than the zero-shot setup discussed above. Nevertheless, both Google and FSMT did not yield scores comparable to monolingual setup. Besides, surprisingly Google yielded worse results than FSMT.

\section{Results \& Discussion}

Table \ref{tab:table_2} shows the best scores of both multilingual and cross-lingual experiments. In multilingual setup mBART performs better than baselines and mT$5$ for both English and Russian. Note that the table shows only the best results of the models. It is also notable that for mT$5$ increased training size for English data provides better metrics for English while keeping metrics for Russian almost the same. We also depict some of the generated detoxified sentences in the Table \ref{tab:table_2} in the part \ref{sec:appendix_b} of Appendix.

As for cross-lingual style transfer, results are negative. None of the models have coped with the task of cross-lingual Textual Style Transfer. That means that models produce the same or almost the same sentences for the language on which they were not fine-tuned so that toxicity is not eliminated. We provide only some scores here in the Table \ref{tab:results} for reference. 

Despite the fact that our hypothesis about the possibility of cross-language detoxification was not confirmed, the presence of multilingual models pre-trained in many languages gives every reason to believe that even with a small amount of parallel data, training models for detoxification is possible.

A recent work by \cite{multilingual_tst} shows that cross-lingual formality Textual Style Transfer is possible. \newcite{multilingual_tst} achieve this on XFORMAL dataset \cite{briakou2021xformal} by adding language-specific adapters in the vanilla mBART architecture \cite{liu2020multilingual} - two feed-forward layers with residual connection and layer normalization \cite{bapna-firat-2019-simple, DBLP:conf/icml/HoulsbyGJMLGAG19}. 

We follow the original training procedure described by \newcite{multilingual_tst} by training adapters for English and Russian separately on $5$ million sentences from News Crawl dataset\footnote{\url{https://data.statmt.org/news-crawl/}}. We use batch size of $16$ and $200$ thousand training iterations. We also then train cross-attentions on our parallel detoxifcation data in the same way. However, models tend to duplicate input text without any detoxification. 
Thus, while the exact same original setup did not work for detoxification, more parameter search and optimization could lead to more acceptable results and we consider the approach by \newcite{multilingual_tst} as a promising direction of a future work on multilingual and cross-lingual detoxification. 



\section{Conclusion}

In this work we have tested the hypothesis that multilingual language models are capable of performing cross-lingual and multilingual detoxification. In the multilingual setup we experimentally show that reformulating detoxification (Textual Style Transfer) as a NMT task boosts performance of the models given enough parallel data for training. We beat simple (Delete method) and more strong (condBERT) baselines in a number of metrics. Based on our experiments, we can assume that it is possible to fine-tune multilingual models in any of the $100$ languages in which they were originally trained. This opens up great opportunities for detoxification in unpopular languages.

However, our hypothesis that multilingual language models are capable of cross-lingual detoxification was proven to be false. We suggest that the reason for this is not a lack of data, but the model's inability to capture the pattern between toxic and non-toxic text and transfer it to another language by itself. This means that the problem of cross-lingual textual style transfer is still open and needs more investigation.





\section*{Acknowledgements}

This work was supported by MTS-Skoltech laboratory on AI.

\bibliography{anthology, custom}
\bibliographystyle{acl_natbib}

\newpage
\appendix
\onecolumn
\section{Data}
\subsection{English Dataset}\label{sec:appendix_1}

Table \ref{tab:ParaDetox} shows examples of sentence pairs from ParaDatex parallel detoxification corpora. There are several polite paraphrases for each toxic sentence in this dataset \cite{dementieva2021crowdsourcing}, this is a consequence of the way these parallel data are collected. Leaving only one paraphrase for one source sentence we could get $6000$ unique pairs of toxic sentences and their polite paraphrases. However, in this work we use data as is. 
\medskip

\begin{table*}[h!]
\centering
\begin{tabular}{l|l}
\hline
\color{red} Original & \color{red} my computer is broken and my phone too!! wtf is this devil sh*t???\\ \hline
Detoxed &  My computer is broken and my phone too! So disappointed!\\
& My computer is broken and my phone too, what is this?\\
& Both my computer and phone are broken. \\ \hline
\color{red} Original &  \color{red}sh*t is crazy around here \\ \hline
Detoxed & It is crazy around here. \\ 
& Stuff is crazy around here. \\
& Something is crazy around here.\\ \hline
\color{red} Original & \color{red} delete the page and shut up \\ \hline
Detoxed & Delete the page and stay silent. \\ 
& Please delete the page. \\
& Delete the page.\\ \hline
\color{red} Original & \color{red} massive and sustained public pressure is the only way to get these
b*stards to act. \\ \hline
Detoxed & Massive and sustained public pressure is the only way to get them to act. \\ 
& Massive and sustained pressure is the only way to get these people to act. \\ \hline
\color{red} Original & \color{red} f*ck you taking credit for some sh*t i wanted to do \\ \hline
Detoxed & You are taking credit for something I wanted to do \\ 
& You’re taking credit fro something i wanted to do. \\ \hline
\color{red} Original & \color{red} you gotta admit that was f*ckin hilarious though! \\ \hline
Detoxed & you got to admit that was very hilarious though! \\
& you gotta admit that was hilarious though! \\
\end{tabular}
\caption{Example sentences from ParaDetox parallel detoxification corpora. Sentence in red is original (toxic) sentence, below are its polite paraphrases. Note that for the purpose of an overall correctness explicit words are masked with "*".}
\label{tab:ParaDetox}
\end{table*}
\newpage
\subsection{Russian Dataset}\label{sec:appendix_2}

Table \ref{tab:RuDetox} shows examples from Russian parallel detoxification corpus. 

\begin{table*}[h!]
\centering
\begin{tabular}{l|l}
\hline
\color{red} Original & \color{red} \foreignlanguage{russian}{Х*рню всякую пишут,из-за этого лайка.v Долбо**изм.}\\ \textit{Translation} & \textit{They write all sorts of bullshit, because of this like. Stupidity.}\\ \hline
Detoxed &  \foreignlanguage{russian}{Чушь всякую пишут, из- за этого лайка.}\\
\textit{Translation} & \textit{They write all sorts of nonsense, because of this like.} \\ \hline
\color{red} Original & \color{red} \foreignlanguage{russian}{А нахрена тогда ты здесь это писал?}\\ \textit{Translation} & \textit{Why the f*ck did you post it here?}\\ \hline
Detoxed &  \foreignlanguage{russian}{Зачем ты это писал?}\\
\textit{Translation} & \textit{Why did you post it?} \\ \hline
\color{red} Original & \color{red} \foreignlanguage{russian}{Е*анутые. Отобрать оружие и лодку и штраф тысяч 500}\\ \textit{Translation} & \textit{F*ckers. Take away weapons, boat and give a fine of 500 thousand.}\\ \hline
Detoxed &  \foreignlanguage{russian}{Отобрать оружие и лодку и штраф тысяч 500.}\\
\textit{Translation} & \textit{Take away weapons, boat and give a fine of 500 thousand.} \\ \hline
\color{red} Original & \color{red} \foreignlanguage{russian}{Не поверишь, сколько е**нутых на планете.}\\ \textit{Translation} & \textit{You won't believe how many f*cked up people are on the planet.}\\ \hline
Detoxed &  \foreignlanguage{russian}{Не поверишь сколько таких на планете.}\\
\textit{Translation} & \textit{You won't believe how many people like that are there on the planet.} \\ \hline
\end{tabular}
\caption{Example sentences from Russian parallel detoxification corpora. Sentence in red is original (toxic) sentence, below are its polite paraphrases.}
\label{tab:RuDetox}
\end{table*}

\section{Generation Examples}\label{sec:appendix_b}

Table \ref{tab:results} contains detoxification examples for different models. It is notable that in some cases models generate almost the same results. This can be explained by the similarity of the training procedure and the fact that the reference answer was the same. 

\begin{table*}[h!]
\centering
\begin{tabular}{l|l}
\hline
\color{gray} Original & \color{gray} After all its hard to get a job if you're stupid.
\\ mBART $5000$ ENG & After all its hard to get a job if you don't understand.\\ 
mBART $3000$ ENG & After all its hard to get a job if you are not smart. \\ 
mT5 base (multilingual) & It's hard to get a job if you don't know. \\ \hline
\color{gray} Original &  \color{gray} Fuck minimum security, put him in real prison.\\
mBART $5000$ ENG & Put him in a real prison with minimum security. \\
mBART $3000$ ENG & I don't care about minimum security, put him in real prison. \\
mT5 base & I don't care about minimum security, put him in real prison. \\ \hline
\color{gray} Original & \color{gray} Sure he should get punished ,but three f*cking years?\\
mBART $5000$ ENG & Sure he should get punished, but three years?\\
mBART $3000$ ENG & He should get punished, but three years? \\
mT5 base & He'd get punished, but three years? \\ \hline
Paraphrase &  \foreignlanguage{russian}{Что у этих людей в башке!? походу насрато!}\\
\textit{Translation} & \textit{What's going on in these people's heads!? It looks like shit!} \\ \hline
mBART $5000$ ENG & \foreignlanguage{russian}{Что у этих людей в голове!? Походу бесчеловечно.}\\
mBART $3000$ ENG & \foreignlanguage{russian}{Что у этих людей в голове? Походу ненормально!} \\
mT5 base & \foreignlanguage{russian}{походу этих людей!? походу!} \\ \hline
\end{tabular}
\caption{Some detoxified sentences produced by our fine-tuned models. \textcolor{gray}{Gray text} refers to the original sentence, below are its paraphrases.}
\label{tab:results}
\end{table*}

\end{document}